# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

### v4

@@ 3.1 ADPO as a Unified Anchored Projection
    DPO (binary comparison form). For $|S_{x}|=2$ with candidates $\{i,j\}$ and Bradley–Terry soft targets $q(i\mid\{i,j\})=\sigma(\beta_{r}\Delta R)$ where $\Delta R=R_{i}-R_{j}$ is the reward difference, Eq. (5) is equivalent to DPO up to temperature scaling and reference shift. Specifically, the Bayes-optimal condition $\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))=q_{ij}$ yields $\Delta_{\theta}-\Delta_{\mathrm{ref}}=(\beta_{r}/\beta)\Delta R$. When $\beta=\beta_{r}$ and $\pi_{\mathrm{ref}}$ is uniform (so $\Delta_{\mathrm{ref}}=0$), this is exactly identical to the original DPO objective; otherwise, it differs by a temperature rescaling $\beta_{r}/\beta$ and an anchor-induced shift $\Delta_{\mathrm{ref}}$.
Table 2: ADPO as a unifying projection framework. Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$ recover major learning paradigms. Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((s-s^{\mathrm{ref}})/\tau)$ to the target $q$; the corresponding student policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$ when the optimum is reachable (Prop. 3.2).
Task Target $q(\cdot\mid x)$ Anchor $\pi_{\mathrm{ref}}$ $\tau$ Recovers ($\tilde{p}_{\theta}$ / $\pi_{\theta}$)
SFT data (one-hot/soft) uniform 11 cross-entropy / same
KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}\!\big(q_{T},\mathrm{softmax}(s/T)\big)$ / train at $T$, infer T=1T{=}1
KD (anchored) teacher $\pi_{T}$ $\pi_{T}$ any $\mathrm{KL}\!\big(\pi_{T}\|\mathrm{softmax}((s-s_{T})/\tau)\big)$ / trust region
[…]
DPO (binary) Bradley–Terry / hard uniform any original DPO (pairwise)
ADPO (listwise) Plackett–Luce / soft any (teacher/old) any this work
@@
$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau}.$ (6)
@@
$\tilde{p}_{\theta}(a)\;\propto\;\exp\!\Big(\frac{\log\pi_{\theta}(a)-\log\pi_{\mathrm{ref}}(a)}{\tau}\Big)\;=\;\Big(\frac{\pi_{\theta}(a)}{\pi_{\mathrm{ref}}(a)}\Big)^{1/\tau}.$
@@
$\pi_{\theta}(a)\;\propto\;\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau},$
which requires: (i) overlapping support: $q(a)>0\Rightarrow\pi_{\mathrm{ref}}(a)>0$ almost everywhere; (ii) integrability: $\int\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau}\,da<\infty$ to ensure normalizability. Under these conditions, the solution is unique up to the normalization constant (equivalently, the Radon–Nikodym derivative $d\pi_{\theta}/d\mu$ is determined by the multiplicative form, where $\mu$ is the base measure).
Proof (discrete): $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau\big)=q$ if and only if $(s-s^{\mathrm{ref}})/\tau=\log q+c\mathbf{1}$ for some constant $c$ (reflecting the shift-invariance of softmax), which rearranges to $\pi_{\theta}\propto\pi_{\mathrm{ref}}\,q^{\tau}$. The constant $c$ is absorbed into the proportionality constant. ∎
@@
Listwise subset constraints and convexity. When $S_{x}$ is only a subset (pairwise/listwise ranking), (6) holds within each group, but the global $\pi_{\theta}$ must simultaneously satisfy multiple group constraints. Crucially, $\mathcal{L}_{\mathrm{ADPO}}$ is convex with respect to the anchored logits $u=(s-s^{\mathrm{ref}})/\tau$ within each group (as shown in Theorem 3.6), but the optimization over shared parameters $\theta$ is typically non-convex. The group-wise closed-form solution provides a target, but there may not exist a parameter $\theta$ that simultaneously achieves $\tilde{p}_{\theta}(\cdot|S_{x})=q(\cdot|S_{x})$ for all groups (except in the tabular case or with unlimited expressivity). Thus, in practice, ADPO performs a parameterized fitting that minimizes the expected group-wise KL, balancing the constraints across all training groups.
@@
    Support overlap (discrete): For discrete action/candidate spaces, we assume $q(i|x)>0\Rightarrow\pi_{\mathrm{ref}}(i|x)>0$ for all $i\in S_{x}$, ensuring that anchored logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ are well-defined.
@@
    Expressivity: Closed-form optimal solutions (Proposition 3.2) assume the policy parametrization $\pi_{\theta}$ has sufficient expressivity to represent the target multiplicative form $\pi_{\mathrm{ref}}\cdot q^{\tau}$. In practice (e.g., shared-parameter networks, autoregressive models), we perform an approximate projection in the parametric family.
@@ 3.2 Pairwise Instantiation
where $\beta>0$ is student temperature, $\Delta_{\theta}=s_{i}-s_{j}$, and $\Delta_{\text{ref}}=s_{i}^{\text{ref}}-s_{j}^{\text{ref}}$ with $s_{i}=\log\pi_{\theta}(\tau_{i}|x)$.
@@ 3.3 Listwise Instantiation
Notation for this subsection: $\tau$ = anchor temperature, $\beta_{r}$ = teacher inverse temperature for forming $q$ from rewards, $\tilde{R}_{i}$ = observed rewards, $\hat{R}_{i}$ = transformed rewards.
For group $S_{x}=\{\tau_{1},\ldots,\tau_{P}\}$, the ADPO listwise loss is:
$\mathcal{L}_{\text{group}}^{\text{ref}}=\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})\right],$ (8)
@@
$\tilde{p}_{\theta}(i|S_{x})=\frac{\exp\left((s_{i}-s_{i}^{\text{ref}})/\tau\right)}{\sum_{j\in S_{x}}\exp\left((s_{j}-s_{j}^{\text{ref}})/\tau\right)},$ (9)
and teacher target $q(i|S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$ uses transformed rewards $\hat{R}_{i}$ via: (a) raw: $\hat{R}_{i}=\tilde{R}_{i}$; (b) rank-based Gaussian transform; or (c) KDE-CDF-Logit: $\hat{R}_{i}=\mathrm{logit}(\hat{F}(\tilde{R}_{i}))$ where $\hat{F}$ is the KDE-estimated CDF.
@@
$\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})=-\frac{1}{\tau}\sum_{i\in S_{x}}q(i|S_{x})\big(s_{i}-s_{i}^{\mathrm{ref}}\big)+\log\!\sum_{j\in S_{x}}\exp\!\left(\frac{s_{j}-s_{j}^{\mathrm{ref}}}{\tau}\right).$ (10)
This decomposition makes explicit that the anchored listwise cross-entropy consists of a linear matching term $\langle q,(s-s^{\mathrm{ref}})/\tau\rangle$ and a log-sum-exp normalization. It directly reveals the connection to the convex conjugate form of the softmax, underlying the implicit trust-region regularization discussed below. Since $\mathrm{KL}(q\|\tilde{p}_{\theta})=\mathcal{L}_{\text{group}}^{\text{ref}}+H(q)$ and $H(q)$ is parameter-independent, minimizing (10) is equivalent to minimizing the KL divergence from the teacher distribution $q$ to the student’s anchored distribution $\tilde{p}_{\theta}$.
3.4 Fisher Geometry and Implicit Trust Region
We now formalize how anchoring provides an implicit trust region through Riemannian geometry.
Lemma 3.3 (Implicit trust region; quadratic form at $p=q$).
Let $u=(s-s^{\mathrm{ref}})/\tau$ and $\tilde{p}(u)=\mathrm{softmax}(u)$. Let $u^{\star}=\log q+c\mathbf{1}$ so that $\tilde{p}(u^{\star})=q$ (the softmax is shift-invariant). Define $\delta=u-u^{\star}$ and center it by $q$ as $\bar{\delta}=\delta-\langle q,\delta\rangle\mathbf{1}$. Then, as $u\to u^{\star}$,
$\mathrm{KL}\!\big(q\,\|\,\tilde{p}(u)\big)\;=\;\tfrac{1}{2}\,\bar{\delta}^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2})\;=\;\frac{1}{2\tau^{2}}\,\mathrm{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\tau\log q_{i}\big]\;+\;o(\cdot),$
where the variance is taken with respect to $q$, measuring the deviation of the anchored log-odds from the target $q$. The quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$, which has a null-space along $\mathbf{1}$ (reflecting shift-invariance).
@@ 3.4 Fisher Geometry and Implicit Trust Region
Let $A(u)=\log\sum_{j}e^{u_{j}}$ so that $\mathcal{L}(u)=A(u)-\langle q,u\rangle$. Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form. The centering removes the null-space along $\mathbf{1}$, reflecting the invariance of softmax to additive shifts. ∎
Corollary 3.4 (Local bounds).
Let $\mu(q)>0$ denote the smallest nonzero eigenvalue of $\mathrm{Diag}(q)-qq^{\top}$ (the null-space is along $\mathbf{1}$). Then locally around $u^{\star}$,
@@
where $\bar{\delta}=\delta-\langle q,\delta\rangle\mathbf{1}$ is the $q$-centered deviation. By Pinsker’s inequality, $\|\tilde{p}(u)-q\|_{1}\leq\sqrt{2\mathrm{KL}(q\|\tilde{p}(u))}$, providing a distributional deviation bound.
On the smallest nonzero eigenvalue: For $K=2$, $\mu(q)=2q(1-q)$. For $K>2$, $\mu(q)$ depends on the full spectrum of $q$; a uniform, distribution-independent lower bound is generally loose. When $\min_{i}q_{i}\geq\varepsilon>0$, the smallest nonzero eigenvalue satisfies $\mu(q)\geq\varepsilon$. Proof sketch: For any $v\in\mathbf{1}^{\perp}$, the Rayleigh quotient gives $v^{\top}(\mathrm{Diag}(q)-qq^{\top})v=\sum_{i}q_{i}v_{i}^{2}-(\sum_{i}q_{i}v_{i})^{2}=\mathrm{Var}_{q}[v]\geq\varepsilon\|v\|^{2}$ since $\sum_{i}v_{i}=0$ and the variance of $v$ under any distribution bounded below by $\varepsilon$ is at least $\varepsilon\|v\|^{2}$.
@@
ADPO performs optimization in the anchored $u$-coordinate system: the reference $\pi_{\mathrm{ref}}$ defines the coordinate origin via $u=(s-s^{\mathrm{ref}})/\tau$, but the center of the local quadratic bowl is at $u^{\star}$ satisfying $\tilde{p}_{\theta}(\cdot;u^{\star})=q(\cdot)$, not at the anchor itself. The local curvature at $u^{\star}$ is governed by the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$, which depends only on the target $q$. Thus: the anchor determines the coordinate system, while the target $q$ determines both the minimizer $u^{\star}$ and the local geometry. This contrasts with:
  • Knowledge Distillation (KD): Minimizes $\mathrm{KL}(q\|p_{\theta})$ in the Euclidean probability simplex. While the Fisher curvature exists at any point, there is no fixed reference anchor, so the geometry lacks a stable coordinate origin.
  • TRPO: Constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in parameter space, defining a moving local quadratic region that requires explicit constraint optimization.
  • ADPO: Anchoring logits to $s^{\mathrm{ref}}$ yields a distribution-space analogue of TRPO with a fixed coordinate origin, inducing an implicit trust region without explicit constraints.
Figure 1: ADPO’s anchored trust region. The Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space, expressed in anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau$; the quadratic well is centered at $u^{\star}$ (i.e., at the target $q$), not at the anchor itself. The green trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift. This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
3.5 Geometric Interpretation: KL-Mirror Descent in $u$-Space
Recall $A(u)=\log\!\sum_{j}\exp(u_{j})$ and $\tilde{p}(u)=\nabla A(u)=\mathrm{softmax}(u)$, and our anchored objective $\mathcal{L}(u)=A(u)-\langle q,u\rangle=\mathrm{KL}\!\big(q\;\|\;\tilde{p}(u)\big)+H(q)$. Let $u_{t}$ be the current anchored logits and $g_{t}=\nabla\mathcal{L}(u_{t})=\tilde{p}(u_{t})-q$.
Theorem 3.5 (KL-mirror descent / Bregman-prox step in $u$-space).
If one were to directly optimize $\mathcal{L}$ over the anchored logits $u$ (rather than shared parameters $\theta$), then for any step size $\eta>0$, the one-step update can be written equivalently as the solution of the following proximal subproblem:
$u_{t+1}\ =\ \arg\min_{u}\ \Big\langle g_{t},\ u\Big\rangle\ +\ \frac{1}{\eta}\,\mathrm{KL}\!\Big(\tilde{p}(u_{t})\,\big\|\,\tilde{p}(u)\Big).$ (11)
Equivalently, since $D_{A}(u\|u_{t})=A(u)-A(u_{t})-\langle\nabla A(u_{t}),u-u_{t}\rangle=\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u))$, (11) is a standard mirror descent / Bregman-prox step with generator $A$.

### v5

@@ 3.1 Core Formulation: Anchored Distribution as the Protagonist
    DPO (binary preferences). For $|S_{x}|=2$ with Bradley–Terry targets $q(i\mid\{i,j\})=\sigma(\beta_{r}\Delta R)$, Eq. (5) reduces to DPO [1] when $\beta=\beta_{r}$ and $\pi_{\mathrm{ref}}$ is uniform.
Remark 3.3 (Connection to AWR/AWAC/MPO).
ADPO also connects to offline RL methods at the $\tilde{p}-$distribution level; see Appendix A for details on this partial equivalence.
Table 2: ADPO as a unifying projection framework. Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning paradigms. Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}})$ to the target $q$; the corresponding policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ when the optimum is reachable (Prop. 3.4).
Task Target $q(\cdot\mid x)$ Anchor $\pi_{\mathrm{ref}}$ $\tau_{\mathrm{anc}}$ Recovers
SFT data (one-hot/soft) uniform 11 cross-entropy
KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}(q_{T},\mathrm{softmax}(s/T))$
KD (anchored) teacher $\pi_{T}$ $\pi_{T}$ any teacher-anchored KD
Max-ent RL $\propto\exp(Q/\alpha)$ uniform 11 $\pi_{\theta}\propto\exp(Q/\alpha)$
DPO (binary) Bradley–Terry / hard uniform any original DPO
ADPO (listwise) Plackett–Luce / soft any any this work
@@
$\pi_{\theta}(i\mid x)\;\propto\;\pi_{\mathrm{ref}}(i\mid x)\cdot q(i\mid x)^{\tau_{\mathrm{anc}}}.$ (8)
@@
$\tilde{p}_{\theta}(a)\;\propto\;\exp\!\Big(\frac{\log\pi_{\theta}(a)-\log\pi_{\mathrm{ref}}(a)}{\tau_{\mathrm{anc}}}\Big)\;=\;\Big(\frac{\pi_{\theta}(a)}{\pi_{\mathrm{ref}}(a)}\Big)^{1/\tau_{\mathrm{anc}}}.$
@@
$\pi_{\theta}(a)\;\propto\;\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau_{\mathrm{anc}}},$
which requires: (i) overlapping support: $q(a)>0\Rightarrow\pi_{\mathrm{ref}}(a)>0$ almost everywhere; (ii) integrability: $\int\pi_{\mathrm{ref}}(a)\cdot q(a)^{\tau_{\mathrm{anc}}}\,da<\infty$ to ensure normalizability. Under these conditions, the solution is unique up to the normalization constant (equivalently, the Radon–Nikodym derivative $d\pi_{\theta}/d\mu$ is determined by the multiplicative form, where $\mu$ is the base measure).
Proof (discrete): $\tilde{p}_{\theta}=\mathrm{softmax}\big((s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=q$ if and only if $(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}=\log q+c\mathbf{1}$ for some constant $c$ (reflecting the shift-invariance of softmax), which rearranges to $\pi_{\theta}\propto\pi_{\mathrm{ref}}\,q^{\tau_{\mathrm{anc}}}$. The constant $c$ is absorbed into the proportionality constant. ∎
@@
Listwise subset constraints and convexity. When $S_{x}$ is only a subset (pairwise/listwise ranking), (8) holds within each group, but the global $\pi_{\theta}$ must simultaneously satisfy multiple group constraints. Crucially, $\mathcal{L}_{\mathrm{ADPO}}$ is convex with respect to the anchored logits $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ within each group (as shown in Theorem 3.8), but the optimization over shared parameters $\theta$ is typically non-convex. The group-wise closed-form solution provides a target, but there may not exist a parameter $\theta$ that simultaneously achieves $\tilde{p}_{\theta}(\cdot\mid S_{x})=q(\cdot\mid S_{x})$ for all groups (except in the tabular case or with unlimited expressivity). Thus, in practice, ADPO performs a parameterized fitting that minimizes the expected group-wise KL, balancing the constraints across all training groups.
@@
Mirror descent = KL proximal step.
If one directly optimizes over the anchored logits $u$ (e.g., tabular setting), each gradient step is equivalent to a KL-trust-region proximal problem (Theorem 3.7):
$u_{t+1}=\arg\min_{u}\ \langle\nabla\mathcal{L}(u_{t}),\,u\rangle+\tfrac{1}{\eta}\mathrm{KL}(\tilde{p}(u_{t})\|\tilde{p}(u)),$
whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$. This exhibits ADPO as a Bregman-proximal method in distribution space.
Dynamic anchoring: monotonicity within fixed-anchor phases.
When the anchor $\pi_{\mathrm{ref}}$ is held fixed, the loss $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ (where $A(u)=\log\sum_{j}e^{u_{j}}$) is strictly convex on the subspace $\mathbf{1}^{\perp}$. Under either (i) local analysis near $u^{\star}$ or (ii) bounded-domain constraints $\tilde{p}_{\theta}\in[\varepsilon,1-\varepsilon]^{K}$, $\mathcal{L}$ is $\mu$-strongly convex on $\mathbf{1}^{\perp}$, guaranteeing linear convergence $\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\leq(1-\mu\eta)(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star}))$ within each fixed-anchor phase (Theorem 3.8). When the anchor updates ($\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$), the coordinate system changes, and we do not claim monotonicity across anchor updates—only within each phase.
@@
    Support overlap (discrete): For discrete action/candidate spaces, we assume $q(i\mid x)>0\Rightarrow\pi_{\mathrm{ref}}(i\mid x)>0$ for all $i\in S_{x}$, ensuring that anchored logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ are well-defined.
@@
    Expressivity: Closed-form optimal solutions (Proposition 3.4) assume the policy parametrization $\pi_{\theta}$ has sufficient expressivity to represent the target multiplicative form $\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$. In practice (e.g., shared-parameter networks, autoregressive models), we perform an approximate projection in the parametric family.
@@ 3.2 Pairwise Instantiation
where $\beta>0$ is student temperature, $\Delta_{\theta}=s_{i}-s_{j}$, and $\Delta_{\text{ref}}=s_{i}^{\text{ref}}-s_{j}^{\text{ref}}$ with $s_{i}=\log\pi_{\theta}(y_{i}|x)$.
@@
Temperature alignment: The pairwise formulation uses inverse temperature $\beta$, which relates to the listwise anchoring temperature via $\beta=1/\tau_{\mathrm{anc}}$. Specifically, $\tilde{p}_{\theta}(i)=\sigma\!\big((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}\big)=\sigma\!\big(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}})\big)$.
@@ 3.3 Listwise Instantiation
Notation for this subsection: $\tau_{\mathrm{anc}}$ = anchor temperature, $\beta_{r}$ = teacher inverse temperature for forming $q$ from rewards, $\tilde{R}_{i}$ = observed rewards, $\hat{R}_{i}$ = transformed rewards.
For group $S_{x}=\{y_{1},\ldots,y_{P}\}$, the ADPO listwise loss is:
$\mathcal{L}_{\text{group}}^{\text{ref}}=\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i\mid S_{x})\log\tilde{p}_{\theta}(i\mid S_{x})\right],$ (10)
@@
$\tilde{p}_{\theta}(i\mid S_{x})=\frac{\exp\left((s_{i}-s_{i}^{\text{ref}})/\tau_{\mathrm{anc}}\right)}{\sum_{j\in S_{x}}\exp\left((s_{j}-s_{j}^{\text{ref}})/\tau_{\mathrm{anc}}\right)},$ (11)
and teacher target $q(i\mid S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$ uses transformed rewards $\hat{R}_{i}$ via: (a) raw: $\hat{R}_{i}=\tilde{R}_{i}$; (b) rank-based Gaussian transform; or (c) KDE-CDF-Logit: $\hat{R}_{i}=\mathrm{logit}(\hat{F}(\tilde{R}_{i}))$ where $\hat{F}$ is the KDE-estimated CDF.
@@
$\sum_{i\in S_{x}}q(i\mid S_{x})\log\tilde{p}_{\theta}(i\mid S_{x})=-\frac{1}{\tau_{\mathrm{anc}}}\sum_{i\in S_{x}}q(i\mid S_{x})\big(s_{i}-s_{i}^{\mathrm{ref}}\big)+\log\!\sum_{j\in S_{x}}\exp\!\left(\frac{s_{j}-s_{j}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\right).$ (12)
This decomposition makes explicit that the anchored listwise cross-entropy consists of a linear matching term $\langle q,(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}\rangle$ and a log-sum-exp normalization. It directly reveals the connection to the convex conjugate form of the softmax, underlying the implicit trust-region regularization discussed below. Since $\mathrm{KL}(q\|\tilde{p}_{\theta})=\mathcal{L}_{\text{group}}^{\text{ref}}+H(q)$ and $H(q)$ is parameter-independent, minimizing (12) is equivalent to minimizing the KL divergence from the teacher distribution $q$ to the student’s anchored distribution $\tilde{p}_{\theta}$.
3.4 Formal Statements of Core Results
We now provide formal statements and proofs for the results sketched in Section 3.
Lemma 3.5 (Implicit trust region).
Let $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ and $\tilde{p}(u)=\mathrm{softmax}(u)$. Let $u^{\star}=\log q+c\mathbf{1}$ so that $\tilde{p}(u^{\star})=q$. Define $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$, the $q$-centered deviation. Then, as $u\to u^{\star}$,
$\mathrm{KL}\!\big(q\,\|\,\tilde{p}(u)\big)\;=\;\tfrac{1}{2}\,\bar{\delta}^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2}).$
The curvature is given by the softmax Fisher metric $\operatorname{Diag}(q)-qq^{\top}$, which is independent of the anchor choice and has a null-space along $\mathbf{1}$.
@@ 3.4 Formal Statements of Core Results
Let $A(u)=\log\sum_{j}e^{u_{j}}$ so that $\mathcal{L}(u)=A(u)-\langle q,u\rangle$. Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form. The centering removes the null-space along $\mathbf{1}$, reflecting the invariance of softmax to additive shifts. ∎
Corollary 3.6 (Spectral bounds).
Let $\mu(q)>0$ denote the smallest nonzero eigenvalue of $\operatorname{Diag}(q)-qq^{\top}$. In the quotient space $\mathbb{R}^{K}/\mathrm{span}\{\mathbf{1}\}$, the objective is strongly convex. Locally around $u^{\star}$,
@@
where $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$. By Pinsker’s inequality, $\|\tilde{p}(u)-q\|_{1}\leq\sqrt{2\mathrm{KL}(q\|\tilde{p}(u))}$.
For $K=2$: $\mu(q)=2q(1-q)$. For $K>2$, when $\min_{i}q_{i}\geq\varepsilon>0$, there exists a constant $c(\varepsilon,K)>0$ such that $\mu(q)\geq c(\varepsilon,K)$ on the quotient space. In particular, for any $v$ satisfying $\langle q,v\rangle=0$,
$\operatorname{Var}_{q}[v]=\sum_{i}q_{i}v_{i}^{2}\geq\varepsilon\|v\|^{2}.$
@@
The anchor $\pi_{\mathrm{ref}}$ defines the coordinate origin, while the target $q$ determines the minimizer $u^{\star}$ and local curvature. This provides a distribution-space analogue of TRPO but with an implicit (Fisher-metric) trust region—no explicit KL constraint needed.
Figure 1: ADPO’s anchored trust region. The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ induces a quadratic “bowl” in the distribution space, expressed in anchored coordinates $u=(s-s^{\mathrm{ref}})/\tau_{\mathrm{anc}}$; the quadratic well is centered at $u^{\star}$ (i.e., at the target $q$), not at the anchor itself. The green trajectory shows the optimization path descending toward the optimal policy, with the anchored coordinate system providing an implicit trust region that prevents large distributional drift. This geometric structure holds for both fixed anchors (stable centers for distillation) and dynamic anchors (adaptive regions for exploration).
3.5 KL-Mirror Descent Interpretation
Theorem 3.7 (KL-mirror descent / Bregman-prox step).
If one directly optimizes $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ over anchored logits $u$ (where $A(u)=\log\sum_{j}e^{u_{j}}$), then for any step size $\eta>0$, the one-step update is equivalent to:
$u_{t+1}\ =\ \arg\min_{u}\ \Big\langle\nabla\mathcal{L}(u_{t}),\ u\Big\rangle\ +\ \frac{1}{\eta}\,\mathrm{KL}\!\Big(\tilde{p}(u_{t})\,\big\|\,\tilde{p}(u)\Big),$ (13)
whose solution is the multiplicative-weights update $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$.